# MoLAN: A Unified Modality-Aware Noise Dynamic Editing Framework for Multimodal Sentiment Analysis

### v1

@@ Introduction
where $\mathcal{L}_{\text{task}}$ involves regression and classification tasks. For regression tasks, we adopt the L1 loss. For classification tasks, the standard cross-entropy loss is employed to optimize the model. The loss of the predicted value $\hat{y}$ and the ground truth $y$ is:
$\mathcal{L}_{\text{task}}=\begin{cases}\mathcal{L}_{\text{reg}}=\frac{1}{N}\sum\limits_{i=1}^{N}|y_{i}-\hat{y}_{i}|&\text{(Regression)}\\ \mathcal{L}_{\text{cla}}=\frac{1}{N}\sum\limits_{i=1}^{N}-y_{i}\log(\hat{y}_{i})&\text{(Classification)}\end{cases}$ (17)
@@
Experiments
Datasets and Evaluation Metrics
We conduct experiments on CMU-MOSI (Zadeh et al. 2016), CMU-MOSEI (Zadeh et al. 2018a), CH-SIMS (Yu et al. 2020), and IEMOCAP (Busso et al. 2008). For the CMU-MOSI and CMU-MOSEI datasets, we follow prior works (Wu et al. 2024; Jiang et al. 2024; Zhuang et al. 2024) to evaluate both regression and classification tasks. For regression, we report the Mean Absolute Error (MAE) and Correlation coefficient (Corr). For classification, we calculate the Acc2 and F1 scores for both the including zero sentiment scores as positive (ACC2Has0 / F1Has0) and the ignoring zero sentiment scores (ACC2Non0 / F1Non0). Additionally, we report ACC7. For the CH-SIMS dataset, we adopt ACC2 and F1 scores as evaluation metrics for the classification task. For the IEMOCAP dataset, we use Weighted-F1 and Macro-F1 scores to assess task performance.
Baselines and Implementation Detail
In the framework experiments, we select influential and reproducible multimodal sentiment analysis models as comparative baselines: MulT (Tsai et al. 2019a), SPECTRA (Yu et al. 2023), KuDA (Feng et al. 2024), SFTTR (Sun and Tian 2025b), MMML (Wu et al. 2024).
In the methodological experiments, in addition to the baseline models used in the framework experiments, we also compare a series of representative approaches for multimodal sentiment analysis: TFN (Zadeh et al. 2017), LMF (Liu et al. 2018), MFM (Tsai et al. 2019b), Self-MM (Yu et al. 2021), UniMSE (Hu et al. 2022), CHFN (Guo et al. 2022), ALMT (Zhang et al. 2023), EMT (Sun et al. 2023), GLoMo (Zhuang et al. 2024), JOSFD (Jiang et al. 2024), t-HNE (Li and Li 2025). t-HNE is the latest MSA denoising method. MMML is the current SOTA MSA model.
@@
Table 2: The performance of the MoLAN framework on the SIMS and IEMOCAP. We perform significance testing on five experimental groups, with p-value of $4.26\times 10^{-2}$, $6.01\times 10^{-4}$, $2.95\times 10^{-3}$, $1.08\times 10^{-3}$, and $1.19\times 10^{-2}$, all of which $<$ 0.05 indicate significant differences.
All experiments are conducted using the PyTorch framework on a hardware setup with 8 RTX A6000 GPUs. For the framework experiments, we strictly follow the hyperparameter settings reported in the original papers to ensure fair and consistent comparisons.
@@
Framework Overall Analysis
Table 3: The performance of the MoLAN+ method on the MOSI and MOSEI. $\lozenge$ indicates our reproduced results, while the other results are reported from the original papers. $\dagger$ indicates MSA denoising method. We perform significance testing on MMML and MoLAN+, with p-value of $7.52\times 10^{-4}$ $<$ 0.05 indicates significant differences. The best results are in bold.
Table 4: Performance of MoLAN+ on the SIMS and IEMOCAP. The significance testing on MMML and MoLAN+, with p-value (0.01 $<$ 0.05) indicating significant differences.
Framework effectiveness. We extensively validate the MoLAN framework using five different methods in four standard datasets, with detailed experimental results shown in Table 1 and Table 2. The results clearly demonstrate that after integrating our framework, the performance of models improves. This phenomenon shows that the MoLAN framework suppresses noise while preserving essential information by performing fine-grained noise editing on modality information, thereby improving the performance upper limit of the model. Notably, even current SOTA method MMML benefits from performance enhancements upon applying our framework. These results further illustrate that existing methods inadequately address denoising.
Framework generalization. Furthermore, we observe that regardless of the specific model to which the framework is applied, the performance consistently exhibits a stable upward trend. This strongly validates the excellent generalization capability of our framework across various models and datasets. The performance of MoLAN on multiple experiments further demonstrates the generalization.
Effectiveness of Dynamic Strategy
@@
Methodological Experiments
We conduct experiments on the MoLAN+ method using four datasets. As shown in Table 3 and Table 4, MoLAN+ outperforms the existing baselines on four datasets and achieves SOTA performance. This indicates that the introduction of the noise-suppression cross attention and noise-driven contrastive learning effectively helps the model capture more representative features, reducing the interference of noise on model performance.
@@
As shown in Figure 4, we conduct a case study comparing the original sota model MMML and the latest denoising model t-HNE. In the example on the left of the figure, we compare it with MMML. It can be seen that MMML pays equal attention to modality information in different regions/segments and does not eliminate the influence of noise. In the example on the right of the figure, we compare it with t-HNE. It can be seen that t-HNE uses a consistent denoising intensity in denoising, which leads to the loss of necessary information and causes incorrect sentiment analysis.
Ablation Study
Table 5: Ablation studies. MulT+ means MulT+MoLAN. DE indicates that dynamic editing. MB indicates that modality-aware blocking. NC means noise-suppressed cross attention. DC means denoise-driven contrastive learning.
@@
As shown in the upper part of Table 5, we first conduct ablation experiments on two parts of MoLAN framework.
@@
Effectiveness of noise dynamic editing. We apply a uniform denoising strength to the modality information. Experimental results (w/o DE) show that the model performance degrades after removing dynamic editing. This indicates that uniform denoising strength cannot handle the differences in noise levels between different regions.This method under-denoises the high-noise regions and over-denoises the low-noise regions, resulting in information loss.
Effectiveness of modality-aware blocking. In addition, we conduct an ablation study on the modality-aware blocking (w/o MB). We use uniform one-dimensional blocking to all modalities. The experimental results show that the model performance has degraded. This shows that the information distribution between different modalities is different, and only by adopting a targeted blocking strategy can the noise be located and processed more effectively.
@@
As shown in the lower part of Table 5, we perform ablation experiments of the two modules of MoLAN+.
@@
Effectiveness of noise-suppressed cross attention. We remove the noise-suppressed cross attention (w/o NC). The results show a decline in model performance. This suggests that denoising information is crucial for the computation of the cross-attention mechanism. After removing the mask matrix, the model’s ability to perceive noise decreases, leading to an overall performance drop. Therefore, incorporating the mask matrix with denoising information can effectively guide the cross-attention computation.
Effectiveness of denoise-driven contrastive learning. We remove the denoise-driven contrastive learning (w/o DC). The decline of performance indicates that contrastive learning plays a key role in the encoding of modality information during optimization. After removing this component, the model cannot fully utilize the relationship between denoised modality information and initial information, leading to a reduction in its ability to suppress noise during the feature extraction process. Therefore, noise-driven contrastive learning promotes the encoding of modality information. More ablation experiments see supplementary material.
Figure 4: Case Study. Blue indicates the target person.
Conclusion
@@
In this paper, we propose MoLAN, a unified modality-aware noise dynamic editing framework. MoLAN adopts a differentiated block strategy to divide different modalities into blocks. Subsequently, the adaptive denoising strength is dynamically allocated according to the noise level and semantic relevance of each block, thus achieving noise editing. As a unified framework, MoLAN can be flexibly integrated into different MSA methods to improve the performance of the original model. In addition, based on the framework, we propose an MSA method MoLAN+. MoLAN+ introduces a noise suppression cross-attention mechanism and denoising-driven contrastive learning to guide the model to focus on essential information. Experiments on five models and four datasets verify the broad effectiveness of the MoLAN. Extensive evaluation on four multimodal benchmark datasets demonstrates that MoLAN+ achieves state-of-the-art performance. In the future, we will further explore the application potential of MoLAN in other multimodal tasks.
@@ Appendix A Pilot Study
The initial modality feature embedding, as the initial information after encoding, serves as the sole foundation for sentiment analysis by the model. However, existing research generally neglects the denoising of the encoded information from each modality and fails to thoroughly explore the impact of noise in the initial feature embedding. Noise within the initial features may interfere with the sentiment analysis outcomes. Enhancing the quality of initial features through effective denoising techniques holds promise for further improving the model’s performance. Therefore, we design and conduct a series of experiments to systematically analyze the noise present in the initial features.
Inspection Method
To investigate the impact of noise in initial multimodal feature information, we design a random masking strategy. Specifically, according to a predefined masking ratio, we randomly select elements within the feature embedding for masking.
@@
where, $\mathbf{f}$ denote the feature embedding matrix, and $\mathbf{M}$ the corresponding mask matrix. Each element $R_{ij}$ is a random number sampled from a uniform distribution over the interval [0,1][0,1][ 0 , 1 ]. The masking ratio is defined by $p\in[0,1]$.
@@
[…]
Multimodal Noise
We first conduct a holistic modality masking experiment, in which we apply uniform masking to all modalities except for text. The models involved include widely used MulT models (Tsai et al. 2019a) as well as the latest SOTA models MMML (Wu et al. 2024). For visual features and audio features, we apply random masking at the same ratio. As shown in Table 6, the results indicate that as the masking ratio increases, model performance improves. This phenomenon suggests that the removed information may contain more noise than useful content, allowing the model to perform better after eliminating these redundant features. These results further confirm that a large amount of noise exists in the initial modality features, and demonstrate that effective denoising strategies are likely to improve model performance.
Unimodal Noise
@@
As shown in Table 7, the audio and visual modalities exhibit performance trends that differ from those of the text modality. As the masking ratio increases, the performance of the audio and visual modalities shows fluctuating improvements. However, performance consistently declines in the text modality. This observation suggests that the initial text modality features are higher quality, containing less noise. Therefore, the text modality can serve as a reference standard. Moreover, the performance improvements in the audio and visual modalities are triggered at different masking ratios, further validating the different distribution of noise across modalities. The performance under single modality masking surpasses that of simultaneously masking two modalities, indicating that applying a uniform masking strategy across all modalities leads to the loss of non-noise information. Therefore, differentiated denoising strategies should be adopted for different modalities.
Appendix B Block Strategy Selection
Using square roots as the basis for block division can achieve relatively average division. This can avoid uneven noise distribution caused by too large a block size, which in turn causes excessive deletion of effective information or incomplete noise removal. It can also avoid the problem of incomplete semantics caused by too small a block size. As shown in Table 8, we design comparative experiments with various block sizes for visual and audio.
Experimental results show that in visual and audio information processing, too large or too small block size will lead to lower ACC values, while factors close to the square root as block size can fully capture effective information and maintain information integrity. Overall, the square root as a block division basis has shown good results in both visual and audio modality, avoiding potential problems caused by too large or too small block size.
[…]
Appendix C Implementation Detail

### v2

@@ 4.4 Sentiment Prediction
where $\mathcal{L}_{\text{task}}$ involves regression and classification tasks. For regression tasks, we adopt the L1 loss, following prior works Mai et al. (2023, 2020), which measures the absolute difference. For classification tasks, the standard cross-entropy loss is employed to optimize the model. The loss of the predicted value $\hat{y}$ and the ground truth $y$ is:
$\mathcal{L}_{\text{task}}=\begin{cases}\mathcal{L}_{\text{reg}}=\frac{1}{N}\sum\limits_{i=1}^{N}|y_{i}-\hat{y}_{i}|&\text{(Reg.)}\\ \mathcal{L}_{\text{cla}}=\frac{1}{N}\sum\limits_{i=1}^{N}-y_{i}\log(\hat{y}_{i})&\text{(Cls.)}\end{cases}$ (17)
@@
5 Experiments
@@
Table 1: The performance of the MoLAN framework on the MOSI and MOSEI. The baseline results in the experiment are obtained through replication. Two evaluation metrics, ACC and F1, are adopted, specifically ACC2Has0 / ACC2Non0 and F1Has0 / F1Non0. $\S$ denotes fine-tuning with LoRA. We perform significance testing on seven experimental groups, with p-value of $5.33\times 10^{-5}$, $1.43\times 10^{-6}$, $4.87\times 10^{-10}$, $2.68\times 10^{-6}$, $2.16\times 10^{-3}$, $2.45\times 10^{-5}$, and $2.66\times 10^{-6}$, all of which << 0.05 indicate significant differences.
Table 2: The performance of the MoLAN framework on the SIMS and IEMOCAP. $\S$ denotes fine-tuning with LoRA. Significance testing in Appendix C.
5.1 Datasets and Baselines
We use CMU-MOSI Zadeh et al. (2016), CMU-MOSEI Zadeh et al. (2018a), CH-SIMS Yu et al. (2020), and IEMOCAP Busso et al. (2008).
Framework: We compare our method with MulT Tsai et al. (2019a), SPECTRA Yu et al. (2023), KuDA Feng et al. (2024), SFTTR Sun and Tian (2025a), MMML Wu et al. (2024), as well as MLLM-based baselines LLaVA-NeXT Li et al. (2024a) and Qwen2.5-VL Bai et al. (2025).
Models: We further benchmark against representative MSA methods, including TFN Zadeh et al. (2017), LMF Liu et al. (2018), MFM Tsai et al. (2019b), Self-MM Yu et al. (2021), UniMSE Hu et al. (2022b), CHFN Guo et al. (2022), ALMT Zhang et al. (2023), EMT Sun et al. (2023), GLoMo Zhuang et al. (2024), JOSFD Jiang et al. (2024), t-HNE Li and Li (2025), and MMML Wu et al. (2024). More details are in Appendix B.
5.2 Framework Overall Analysis
This section provides an overall analysis of the proposed framework to assess its performance across different evaluation dimensions. We first examine the framework’s effectiveness in enhancing MSA, followed by an analysis of its universality ability.
Table 3: The performance of the MoLAN+ on the MOSI and MOSEI. $\lozenge$ indicates our reproduced results. $\dagger$ indicates MSA denoising method. $\S$ denotes fine-tuning with LoRA. We perform significance testing on MMML and MoLAN+, with p-value of $7.52\times 10^{-4}$ << 0.05 indicates significant differences. The best results are in bold.
Table 4: Performance of MoLAN+ on the SIMS and IEMOCAP. Significance testing in Appendix C.
Framework effectiveness. We validate the proposed MoLAN framework on four standard MSA datasets, using five representative MSA models and two MLLMs. Detailed experimental results are shown in Table 1 and Table 2. Five MSA models all show dramatically performance improvements after integrating the MoLAN framework. This trend demonstrates that MoLAN can be universally applied across diverse model architectures, effectively improving the quality of multimodal feature fusion. This further demonstrates that MoLAN, through fine-grained noise editing of modality information, can suppress irrelevant noise while retaining key information, thereby strengthening semantic alignment and collaborative representation across modalities. It is worth noting that even the current SOTA, MMML, achieves further improvement after integrating MoLAN. This observation indicates that existing methods still have limitations in modality denoising, and that MoLAN helps overcome these performance bottlenecks. Furthermore, experiments on MLLMs further validate the framework’s effectiveness. Both LLaVA-NeXT and Qwen2.5-VL exhibit consistent performance improvements after integrating MoLAN. Result demonstrates that MoLAN is not only applicable to traditional MSA models but also seamlessly integrates with mainstream MLLM architectures.
Framework universality. Experimental results show that whether MoLAN is integrated into dedicated MSA models or deployed on MLLMs, it consistently exhibits significant performance improvements. This consistent trend verifies the strong universality capability of MoLAN, indicating that it effectively adapts to different architectures, task settings, and data distributions. Overall, the results suggest that MoLAN serves as a general enhancement module for MSA tasks, providing a reliable solution for multimodal representation learning.
@@ 5.2 Framework Overall Analysis
5.3 Effectiveness of Dynamic Strategy
@@
5.4 MoLAN+ Experiments
We conduct a comprehensive evaluation of MoLAN+ on four datasets. As presented in Table 3 and Table 4, MoLAN+ consistently achieves SOTA performance across all datasets, surpassing both existing baseline models and recent MLLM-based approaches. This superior performance demonstrates that the noise-suppression cross attention and noise-driven contrastive learning modules effectively enhance the model’s discriminative capability and semantic alignment across modalities. By emphasizing emotion-relevant multimodal representations during feature extraction, the model is able to mitigate the influence of noisy information and maintain stable performance across diverse conditions. Moreover, when compared with the framework experiments, we observe that although other models benefit from the integration of the MoLAN framework, their results still fall short of the overall performance achieved by MoLAN+. These findings suggest a strong synergy between noise-suppression cross attention, noise-driven contrastive learning, and the MoLAN framework, jointly contributing to the superior adaptability and effectiveness of MoLAN+ in MSA tasks.
@@ 5.4 MoLAN+ Experiments
5.5 Ablation Study
Table 5: Ablation studies.
@@ 5.5 Ablation Study
As shown in the upper part of Table 5, we first conduct ablation experiments on two parts of MoLAN.
@@
Effectiveness of noise dynamic editing. We apply a uniform denoising strength to the modality information. Experimental results (w/o DE) show that the model performance degrades after removing dynamic editing. This indicates that uniform denoising strength cannot handle the differences in noise levels between different regions. Such denoising may result in insufficient denoising in high-noise areas, while low-noise areas may lose valuable information due to over-denoising. The experimental results further verify the necessity of dynamic editing, which adaptively adjusts denoising strength based on local noise characteristics to achieve a better balance between noise suppression and information preservation.
We independently apply the denoising framework to the visual modality (MoLAN(v)) and the audio modality (MoLAN(a)). The results show that although the single-modality denoising does not reach the performance level of joint multimodal denoising, it still achieves a significant improvement compared with the model without the denoising framework. This finding indicates that the denoising mechanism also plays a positive role in single-modality denoising, while the collaborative removal during joint multimodal denoising further enhances the overall performance.
Effectiveness of modality-aware block partitioning. We conduct an ablation study on the modality-aware block partitioning strategy to examine its effectiveness. Specifically, we first apply a uniform one-dimensional block partitioning to all modalities (w/o MB), and then a uniform two-dimensional block partitioning to both the visual and audio modalities (v(2D), a(2D)). The results show a noticeable decline in performance, indicating that the information distribution differs significantly across modalities. These findings suggest that only a targeted block partitioning strategy can effectively localize and suppress modality-specific noise. Furthermore, when we assign one-dimensional block partitioning to the visual modality and two-dimensional block partitioning to the audio modality (v(1D), a(2D)), the model performance also decreases. This further verifies that the adopted configuration, namely 2D block partitioning for visual and 1D block partitioning for audio, better aligns with the properties of each modality and thus enables more efficient denoising.
@@
As shown in the lower part of Table 5, we perform ablation experiments for MoLAN+.
@@
Effectiveness of noise-suppressed cross attention. We remove the noise-suppression cross-attention mechanism (w/o NC) to examine its effect. The experimental results show a performance drop, indicating that denoising information plays a crucial role in cross-attention computation. Specifically, when the mask matrix is removed, the model’s ability to perceive and suppress noise weakens, leading to an overall decline in performance. These results suggest that explicitly incorporating denoising information into the mask matrix effectively guides the cross-attention mechanism to focus on informative regions, reduce noise interference, and enhance the model’s noise resistance.
Effectiveness of denoising-driven contrastive learning. We further remove the denoising-driven contrastive learning module (w/o DC) to validate its contribution. The results reveal a drop in performance, suggesting that this module is essential for learning stable and discriminative modality representations. In particular, the contrastive objective establishes a semantic constraint between denoised and original features, enabling the model to maintain both distinctiveness and consistency of feature distributions under noisy conditions. Once this mechanism is removed, such relational constraints are weakened, making it difficult for the model to separate noise from informative signals in the latent space. Therefore, the denoising-driven contrastive learning not only improves cross-modal alignment but also enhances the model’s universality ability in complex noisy environments. Additional analyses and case studies in Appendix D and Appendix E.
6 Conclusion
@@
We propose MoLAN, a unified modality-aware noise dynamic editing framework that partitions modalities into blocks and dynamically assigns denoising strength based on each block’s noise level and semantic relevance. MoLAN is plug-and-play and can be integrated into various MSA models and MLLMs to improve performance. Built on MoLAN, MoLAN+ further introduces noise suppression cross-attention mechanism and denoising-driven contrastive learning to emphasize essential information. Extensive experiments demonstrate the effectiveness of MoLAN and MoLAN+. Overall, our work provides a practical pathway for enhancing robustness in multimodal systems.
Limitations
Our current evaluation mainly covers several representative MSA benchmarks to demonstrate the generality and integrability of MoLAN and MoLAN+. However, we do not yet conduct systematic evaluations under more challenging settings, such as cross-domain generative tasks or more complex real-world scenarios. In these settings, noise patterns and cross-modal interaction modes can be more diverse, thereby posing new requirements for noise editing. Therefore, comprehensive experiments and analyses on these broader scenarios remain to be further complemented.
@@ Appendix A Pilot Study
Table 6: Multimodal Noise. $\downarrow$} indicates a performance drop compared to the previous row, while $\uparrow$} indicates an improvement. The metric used is ACC2.
The initial modality feature embedding, serving as the encoded representations of each modality, form the foundation of MSA. However, existing studies generally overlook the denoising of these encoded features and fail to investigate the potential impact of noise within the initial feature representations. Noise within the initial features may interfere with the accuracy of MSA and hinder the model’s ability to capture critical information. By enhancing the quality of the initial features through effective denoising, the overall performance of the model can be further improved. Therefore, we design and conduct a series of validation experiments to systematically analyze the presence and influence of noise in the initial modality features.
A.1 Random Masking Strategy
To investigate the impact of noise in the initial multimodal feature representations, we design a random masking strategy. Specifically, we randomly mask a portion of the elements in the feature embeddings according to a predefined masking ratio. In this way, we observe how the model performs under conditions of partial information loss.
@@ A.1 Random Masking Strategy
where $\mathbf{F}$ denote the feature embedding matrix, and $\mathbf{M}$ the corresponding mask matrix. Each element $R_{ij}$ is a random number sampled from a uniform distribution over the interval $[0,1]$. The masking ratio is defined by $p\in[0,1]$.
@@
A.2 Multimodal Noise
We first apply the random masking strategy to the overall modality-level features and conduct experiments on both the visual and audio modalities with the same masking ratio. As shown in Table 6, the results show that the model performance improves as the masking ratio increases. This phenomenon indicates that the removed portions of the features may contain more noise than useful information. Therefore, eliminating these noisy components allows the model to focus on more critical representations, leading to better performance. These findings provide preliminary evidence that a considerable amount of noise exists in the initial modality features and that not all encoded information contributes positively to the task. Consequently, applying an effective denoising strategy improves the quality of modality representations and ultimately enhances the overall performance of MSA task.
A.3 Unimodal Noise
@@
As shown in Table 7, the audio and visual modalities exhibit performance trends that differ from those of the text modality. As the masking ratio increases, the performance of the audio and visual modalities shows fluctuating improvements. However, performance consistently declines in the text modality. This observation suggests that the initial text modality features are higher quality, containing less noise. Therefore, the text modality can serve as a reference standard. Moreover, the performance improvements in the audio and visual modalities are triggered at different masking ratios, further validating the different distribution of noise across modalities. This observation suggests that the differences in noise characteristics between different modalities may lead to the loss of critical information in some modalities when a unified denoising strategy is applied to all modalities. Therefore, differentiated denoising strategies should be adopted for different modalities.
Appendix B More Details
B.1 Dataset and Metrics
These datasets encompass both Chinese and English corpora, incorporating text, audio, and visual modalities across diverse contexts such as monologues, dialogues, and film clips, with annotations covering both discrete and continuous emotional dimensions. For the CMU-MOSI and CMU-MOSEI datasets, we follow prior works Wu et al. (2024); Jiang et al. (2024); Zhuang et al. (2024) to evaluate both regression and classification tasks. For regression, we report the Mean Absolute Error (MAE) and Correlation coefficient (Corr). For classification, we calculate the Acc2 and F1 scores for both the including zero sentiment scores as positive (ACC2Has0 / F1Has0) and the ignoring zero sentiment scores (ACC2Non0 / F1Non0). Additionally, we report ACC7. For the CH-SIMS dataset, we adopt ACC2 and F1 scores as evaluation metrics for the classification task. For the IEMOCAP dataset, we use Weighted-F1 and Macro-F1 scores to assess the performance of the classification task.
B.2 Baselines
Framework: We select influential and reproducible multimodal sentiment analysis models as comparative baselines: MulT Tsai et al. (2019a), SPECTRA Yu et al. (2023), KuDA Feng et al. (2024), SFTTR Sun and Tian (2025a), MMML Wu et al. (2024). These models cover different research paradigms such as multimodal feature alignment, cross-modality attention modeling, and knowledge enhancement. With the rapid rise of the Multimodal Large Language Model (MLLM), unified understanding and reasoning across modalities has become a key direction in multimodal research. To verify the applicability and stability of our framework within this new paradigm, we further introduce LLaVA-NeXTLi et al. (2024a) and Qwen2.5-VLBai et al. (2025) as representative models for comparison, further comprehensively examining the framework’s universality performance.
Models: In addition to the baseline models used in the framework experiments, we also compare a series of representative approaches for MSA: TFN Zadeh et al. (2017): Models intra- and inter-modality dynamics in an end-to-end manner. LMF Liu et al. (2018): Employs low-rank tensors for efficient multimodal fusion, reducing computational complexity. MFM Tsai et al. (2019b): Jointly optimizes generative and discriminative objectives on multimodal data and labels. Self-MM Yu et al. (2021): Designs a self-supervised label generation module to automatically obtain unimodal supervision signals. UniMSE Hu et al. (2022b): Unifies multimodal sentiment analysis and emotion recognition tasks from multiple perspectives. CHFN Guo et al. (2022): Based on a Transformer architecture, it efficiently fuses unaligned multimodal sequences. ALMT Zhang et al. (2023): Introduces an adaptive hyper-modality learning module to suppress irrelevant or conflicting information under the guidance of language. EMT Sun et al. (2023): Enhances model robustness in scenarios with incomplete modalities while maintaining efficiency and performance. GLoMo Zhuang et al. (2024): Integrates local representations from each modality and combines them with global representations to enhance expressive power. JOSFD Jiang et al. (2024): Incorporates fuzzy logic to model both subjective and objective fuzziness in sentiment information. t-HNE Li and Li (2025): Text-guided hierarchical denoiser improves sentiment analysis performance via two-stage denoising and contrastive learning mechanism. t-HNE is the latest MSA denoising method. MMML is the current SOTA MSA model.
B.3 Implementation Detail